How much water should be added to 25 liters of a 14% salt solution to make a 5% salt solution?
Answer by TimothyLamb(4379)

---
0x + 14*25 = 5(x + 25)
14*25 = 5x + 125
5x = 14*25 - 125
5x = 225
x = 225/5
---
answer:
x = 45 liters
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php